OpenStudy (anonymous):

PLEASE HELP ME :( Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7

OpenStudy (badmood):

4x - 3y = 9 -->(2) -8x + 6y = -18 ---------------- 8x - 6y = 18 (result of multiplying by 2) -8x + 6y = -18 ----------------add 0 + 0 = 0 0 = 0 when the answer comes out equal, example : 0 = 0, 2 = 2, 5 = 5, and so forth, the answer is INFINITE SOLUTIONS.